Evaluate ∫y=0^{1}∫x=y^{2√2-y^{2}}y/2√x^{2}+y^{2}dxdy ,by clange the order of integration.


1
Expert's answer
2021-08-30T01:54:13-0400

"∫_{y=0}^{1}∫_{x=y}^{2\sqrt{2-y^{2}}} \frac{y}{2}\sqrt{x^{2}+y^{2}}dxdy\\
∫_{x=0}^{1}∫_{y=x}^{\sqrt{2-\frac{x^2}{4}}} \frac{y}{2}\sqrt{x^{2}+y^{2}}dydx\\
=\int _0^1\left(\frac{\left(3x^2+8\right)^{\frac{3}{2}}}{48}-\frac{\sqrt{2}x^3}{3}\right)dx\\
=\frac{-48\sqrt{3}\ln \left(2\right)+32\sqrt{3}\ln \left(\sqrt{3}+\sqrt{11}\right)+23\sqrt{11}}{192}-\frac{1}{6\sqrt{2}}\\
=0.44671"